OpenStudy (anonymous):

The lines of the equations are parallel, perpendicular, or neither. y = 2x + 3 y = 2x parallel perpendicular neither

OpenStudy (igreen):

Parallel = Same Slope, Different y-intercepts Perpendicular = Opposite Slopes

OpenStudy (igreen):

First off, tell me what the slopes are for y = 2x + 3 and y = 2x. Both of them are in the form of y = mx + b, where 'm' is the slope.

OpenStudy (anonymous):

???

OpenStudy (igreen):

What's the slope of y = 2x + 3?

OpenStudy (igreen):

It's located right before the 'x'.

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yep, and what's the slope of y = 2x?

OpenStudy (anonymous):

2?

OpenStudy (igreen):

Yep! So the slopes are the same, so the lines are parallel.

OpenStudy (anonymous):

What about these? y = -x + 3 y = -x - 3

OpenStudy (anonymous):

??

OpenStudy (igreen):

Ask a new question :)